Infrared radiation has wavelengths ranging
from about 800 nm to 1 mm. What is the
frequency of radiation of wavelength 990 nm?
Answer in units of s−1
.
Determine the frequency of light with a wavelength of 3.637 × 10−7
cm.
Answer in units of Hz.
1
Expert's answer
2019-10-07T06:41:21-0400

1 Hz = 1 s-1

с/λ = ν

1) (2.99*108)/(990*10-9) = 3.03*1014 s-1

2) (2.99*108)/(3.637*10-7) = 8.24*1014 Hz